the specific weight of an object in 3.5 kn /m^3. determine the (a) mass density (b) mass if the dimention of the object are 6 cm by 15 cm by 8 cm (c) specific volume, and (d) specific gravity. solution


1
Expert's answer
2021-03-09T07:43:36-0500

"Mass density= 3.5kN/m^3\times1,000=3,500Kg/m^3"

Mass density=3500


"Mass=density\times volume"


"Volume=length\times width\times height"

"Volume=0.15\times 0.08\times 0.06=0.00072m^3"


"Mass=3,500kg/m^3\times 0.00072m^3=2.52Kg"


Mass = 2.52 Kg


"Specific volume=\frac {volume }{mass}=\frac {0.00072m^3}{2.52kg}=0.0002857m^3/kg"


Specific volume= 0.0002857


"Specific gravity=\frac{Density of oject}{density of water}=\frac{3500kg/m^3}{1000kg/m^3}=3.5"


Specific gravity=3.5